University bookstores order books that instructors adopt for their courses. The number of copies ordered matches the projected demand. However, at the end of the semester the bookstore has to many copies on hand and must return to them to the publisher. A bookstore has a policy that the proportion of books returned should be kept as small as possible. The average is supposed to be less than 10%. To see whether the policy is working, a random sample of 30 book titles was drawn and the fraction of the total originally ordered that are returned is recorded. The sample mean is 8.1% and the sample standard deviation is 3.7%. Find the p-value for this hypothesis test. H_{a} \:: \mu<10\%
Answer #1

Solution :

This is the left tailed test .

t = (0.084 - 0.10) / 0.037 / sqrt 30

= -2.37

Df = 30 - 1 = 29

P value = 0.0123